In this problem you will use variation of parameters to solve the nonhomogeneous equation fy + 4ty + 2y = 1 + 12 A. Plug y
Answer #1

salm: - Giver now-homergen ty +4ty + 2y on- homogeneous equation at² + t² Let Then y = t. = n thi n-2 y and y = n(n-1) thHence in and Ya are linearly independent, a can te written ③ Now equation o 12 y + 2 Y t + 1 t² componing y + pyl + 9 g a Zth Then the general solution is y=+*++*+($.$)++fle = ttht 2 + + £ - - y = Gtt +6E² 4 - 2 2 t 4 t² 20 where G C are constan